[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Oct 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Oct 29 20:38:59 UTC 2017


HSU Shift Log for Sun Oct 29 2017 16:38:53 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues: None.

N-AWIPS issues: None. (radar imagery is being ingested into NAWIPS again)

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171029/6e70082a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list